# In-vitro measurements coupled with in-silico simulations for stochastic calibration and uncertainty quantification of the mechanical response of biological materials

### v1

@@ Abstract
In the standard Bayesian framework, a likelihood function is required, which can be difficult or computationally expensive to derive for non-homogeneous biological materials under complex mechanical loading conditions. Here, we propose a simple and practical approach based on in-vitro measurements and in-silico simulations using the likelihood-free Bayesian inference with the finite element method simultaneously for stochastic calibration and uncertainty quantification of the mechanical response of biological materials. We implement this approach for distal, middle, and proximal human Achilles tendon specimens obtained from diabetic patients post-amputation. A wide range of in-vitro loading conditions are considered, including one-step and two-step relaxation, as well as incremental cyclic loading tests. In-silico simulations are performed for the tendons assuming a fiber-reinforced viscoelastic response, which is modeled separately for the ground matrix and fiber components. Initially, the calibration of the specimen-specific parameters is predicted using Bayesian optimization and the sensitivity of each parameter is evaluated using the Sobol index and random forest. Then, the bounds of these parameters are used as priors, uniformly sampled, and coupled with in-vitro data in simulation-based approximate Bayesian computations to calibrate and quantify the uncertainty parameters for three loading cases. The results demonstrate that increasing the number of loading scenarios improves the accuracy of the specimen-specific best-fit parameter set. Furthermore, approximate Bayesian inference shows that in-silico simulations using the posterior parameters, can capture the uncertainty bounds of in-vitro measurements. This approach provides a useful framework for stochastic calibration of constitutive material model parameters without the need to derive a likelihood function, regardless of the specimen’s geometry or loading conditions.
Keywords: In-vitro measurements, in-silico simulations, biological materials, uncertainty quantification, stochastic material calibration, approximate Bayesian computation, finite element method
@@ 1 Introduction
In-silico mechanical modeling and simulation of biological materials has the potential to facilitate advancements in every aspect of the medical field. It offers numerous advantages, such as a promising tool for pre-clinical testing of medical devices, aiding in the diagnosis and evaluation of targeted treatments for patient-specific applications, and supplementing experimental investigations when in-vivo experiments are not possible [viceconti_possible_2021, smit_silico_2024, VV40, pekedis_2025_2]. It can also model patient-specific tissue or organ-level behavior and, more generally, account for variations within a population. One such factor that significantly influences the in-silico simulation outcome is the constitutive material model, as well as the parameters used in it. The model and its calibrated biomechanical parameters should accurately reflect and mimic the tissue’s behavior. These parameters are typically calibrated using a non-linear least squares fit of the model to mechanical testing data from experiments such as tension [ogden_fitting_2004, pekedis_2021], compression [budday_fifty_2020], and biaxial testing [niestrawska_microstructure_2016]. For most calibration procedures, the choice of a constitutive model is assessed by the quality of fit (e.g., $R^{2}$ and the value of the root mean square error RMSE). This technique is efficient and powerful for patient or specimen-specific calibration. However, because most biological materials are nonhomogeneous and anisotropic, deploying experimental tests using standard mechanical tests may not be enough for mechanical characterization of the tissue. Therefore, calibration of the model parameters directly from the experimental data may not always be possible or relevant. In some studies, the inverse technique is performed to overcome these issues [MENGONI2021105, narayanan_inverse_2021]. In brief, the inverse approach is considered as the estimation of input parameters (material properties, geometric features, loading and boundary conditions, etc.) from output observations (displacement, strains, velocity). Once a forward model is created, the input parameters are predicted using experimental observations with an optimization technique. This method is powerful for patient/specimen-specific material calibration when inter-variability is not considered.
@@
In addition to inhomogeneous characteristics, biological materials often exhibit intrinsic variability in their biomechanical response when subjected to external loads. The inter-subject or inter-specimen variability and uncertainty in biomechanical properties of these materials are high, which may be influenced by many factors such as the donor’s age, genetics, the tissue’s pathological state, gender, and other specific factors related to the tissue characteristics. For in-vitro samples, embalming medium, temperature, humidity and duration are the additional factors that can affect the biomechanical properties. Oftentimes, in-silico model is hindered not only by the incompleteness of the theoretical approach, but also by uncertainties due to mentioned factors that arise in the experimental observations. Thus, there is an urgent need to ensure that it can capture the intrinsic population of variability and not only representative of ‘mean’ individuals.
@@
Uncertainty can also arise due to a lack of knowledge of the physical model of interest (insufficient initial and boundary conditions), as an addendum to the inherent variation in material properties [oberkamf, henninger_validation_2010]. However, most of the studies implemented in the field of biomechanics have mainly investigated uncertainties due to variabilities in material properties using the Bayesian framework [mihai_LA, laz]. In this framework, once the prior distribution of the parameters, in conjunction with a likelihood function, is constructed, the posterior distribution is determined using random sampling methods. There has been much research in the field of brain mechanics [madireddy_bayesian_2015, teferra_bayesian_2019, brewick_uncertainty_2018, staber_stochastic_2017, mihai_LA]. However, the standard Bayesian framework requires a likelihood function, which can be difficult or computationally expensive to derive for nonhomogeneous materials under complex biomechanical loading conditions. To overcome on these such issues, an approximate Bayesian computation (ABC) approach that does not use the likelihood function has been proposed [toni_simulation-based_2010, schalte_pyabc_2022] and implemented recently in some fields including structural health monitoring [Guofeng] medical decision making [Huaqing] and operation research management [garud]. However, the coupling of in-silico simulations for stochastic calibration and uncertainty quantification using likelihood-free Bayesian computation in the field of mechanics of biological materials has been largely unaddressed to date, to the best of our knowledge. The most attractive aspect of this approach is that the likelihood function is not required to be explicitly represented in a closed analytical form. Once an in-silico model is built to mimic the in-vitro conditions, it directly bypasses the need for explicit likelihood evaluation and relies on comparing the simulated data from the in-silico model with experimental observations from the in-vitro test. This approach can be implemented in a wide range of in-vivo and in-vitro applications, ranging from predicting the mechanical characteristics change of tissue due to the risk of pathogenic disease to planning surgical procedures. We note that this study is not targeted to develop or present a constitutive material model for connective tissue, but rather to show how likelihood-free inference can be used directly with the numerical simulator and measurements. Any suggested model can be directly embedded in the simulator.
@@
In this work, we suggest a simple approach based on in-vitro measurements, in-silico simulations, and approximate Bayesian computation for stochastic calibration and uncertainty quantification of material parameters. The paper is structured as follows. First, a specimen-specific calibration procedure using Bayesian optimization is presented. Next, the sensitivity analysis and stochastic calibration steps implemented in ABC are detailed in the following section. Finally, the results obtained from the methodology are presented and summarized, together with some concluding remarks.
@@ 2.1 In-vitro measurements
The approach is performed on recently published data of Achilles tendons from diabetic patients who underwent below or above knee amputation [pekedis_2025]. In summary, the tendons were harvested from the proximal end to the insertion of the Achilles tendon at the calcaneus. Their length was about 120 to 130 cm and were partitioned into three equal sections (distal, middle, and proximal site). Then, they were resectioned to form a standard dog bone specimen. The dimensions of each specimen at the measurement site had a gauge length of approximately 10 mm, a width of $\sim$ 8 mm and a thickness of $\sim$ 4 mm. Uniaxial displacement-type loading was performed to collect force measurements. First, samples were preloaded at 0.1 MPa to achieve parallel alignment of the fibrils, and then preconditioned for 10 cycles with a strain amplitude of 2% strain at a strain rate of 0.5% $s^{-1}$. The test protocols were as follows:
@@
Further details of the tests and in-vitro measurements are reported in previous work [pekedis_2025]. A typical stress-strain observation of a specimen for three loading scenarios is presented in Figure 1. The force and displacement signals have a sampling rate of 100 Hz. In-vitro measurements include 81 datasets (9 specimens x 3 sites x 3 test protocols). Each of these data sets consists of three dimensions: the first is time, the second is displacement, and the third is force. The data sets are organized and saved as PYTHON pickle for further serialization processes between in-vitro measurements and in-silico simulations for sensitivity, optimization and Bayesian inference tasks.
@@ 2.2 Fiber reinforced visco elastic model
Soft biological tissues, such as the Achilles tendon, are collagenous tissues that connect the calf muscle to the calcaneus. Their primary function is to transmit tensile forces and act as mechanical dampers. These tissues are primarily composed of water, Type I collagen fibers, and proteoglycans. They can be viewed as fibrous composites, consisting of a ground matrix embedded with families of collagen fibers. The tendon is modeled using the fiber-reinforced viscoelastic (FRVE) approach, in which the solid matrix is divided into a nonfibrillar part representing the proteoglycan matrix and a fibrillar part representing the collagen fibers. The biomechanical response of these constituents, which drives the approximate Bayesian computation (ABC), is simulated using the FEBio framework [maas_febio_2012]. In this FRVE model, two solid mixtures are considered: the ground matrix, representing the proteoglycan matrix, and the collagen fibrils, representing the fibers. Viscoelastic effects are incorporated for both constituents. Specifically, the ground matrix is modeled as a compressible visco linear elastic material, whereas the collagen fibrils are assumed to be compressible visco nonlinear elastic materials that respond only to tensile loads. The relaxation function for the fibril and matrix components is given by:
@@
where $t$ is time, $g_{i}$ and $\tau_{i}$ are $i$’th viscoelastic characteristic coefficients. Additionally, the strain energy density for the fiber component is follow:
$\psi_{n}=\begin{cases}0<I_{n},&0,\ \text{under compression}\\ \frac{\xi}{\alpha\beta}(\exp[\alpha(I_{n}-1)^{\beta}]-1),&1\leq I_{n}\leq I_{0},\text{for nonlinear toe region}\\ B(I_{n}-I_{0})-E_{f}(I_{n}^{1/2}-I_{0}^{1/2})+\frac{\xi}{\alpha\beta}(\exp[\alpha(I_{0}-1)^{\beta}]-1),&I_{n}\geq I_{o},\ \text{for linear region }\end{cases}$ (2)
where $E_{f}$ is the fiber modulus in the linear region, $\alpha$ is the coefficient of exponential component, $\beta$ is the power-law component in the toe region, $I_{0}=\lambda_{0}^{2}$ and $I_{n}=\lambda_{n}^{2}$. Here, $\lambda_{n}$ represents the current stretch ratio, and $\lambda_{0}$ denotes the stretch ratio in the transition from the toe region to the linear region. Additionally, $\xi$ and $B$ are as follows:
$\xi=\frac{E_{f}(I_{0}-1)^{2-\beta}\exp[-\alpha(I_{0}-1)^{\beta}]}{4I_{0}^{3/2}[\beta-1+\alpha\beta(I_{0}-1)^{\beta}]}$ (3)
$B=E_{f}\frac{2I_{0}[\beta-0.5+\alpha\beta(I_{0}-1)^{\beta}]-1}{4I_{0}^{3/2}[\beta-1+\alpha\beta(I_{0}-1)^{\beta}]}$ (4)
In-vitro measurements showed that $\lambda_{0}$ is approximately 1.01 and is assumed to be a fixed parameter. The above equation indicates that the fiber modulus reduces to zero unless $\beta=2$. Thus, we set $\beta=2$ in the in-silico simulation. In summary, the components constituting the matrix are assumed to exhibit viscoelastic behavior, and the fiber is modeled as a fiber-reinforced viscoelastic material. The descriptions of parameters, the search ranges for Bayesian optimization, and their priors are summarized in Table  1.
Table 1: Parameters used in Bayesian optimization (BO) and approximate Bayesian computation (ABC) for the three sites.
@@
Min and Max values are also used as uniformly distributed prior ranges for ABC.
*represents $\alpha$ search ranges in BO and, **denotes uniformly distributed prior ranges of $\alpha$ for ABC.
@@ 2.3 In-silico simulations
Numerical methods such as the finite element method, meshless methods, or the finite volume method can be used as simulators to bridge the gap between in-vitro measurements and Bayesian computation for stochastic calibration and quantification of uncertainty in the biomechanical response of biological materials. Here, we used the finite element method as a simulator. The geometric dimensions of the in-vitro sample are approximately $10\times 8\times 4\,\text{mm}$ (gauge length $\times$ width $\times$ thickness).
To investigate the effectiveness of the approach and reduce the CPU solution time, we modeled half of the sample by applying a symmetric boundary condition using a single hexagonal element. As stated in the previous section, we considered an FRVE material property for the tendon (Figure 2). The boundary conditions of the in-silico model are similar to the loading conditions of the in-vitro test: the nodes located in the lower section are restricted only in the vertical direction ($u_{z}(t)=0$), and the tensile displacement observed from the in-vitro measurement ($u_{z}(t)$) for the three loading tests is applied longitudinally to the upper four nodes. A symmetric boundary condition is applied to the four nodes on the left side ($u_{y}(t)=0$) (Figure 2). We note that additional in-silico outputs, such as the deformation gradient, the strain distribution, or the direct stress, can also be extracted and used with ABC if the corresponding measurements are available. Typically, full-field deformation measurements can be collected via image-based techniques, such as digital image correlation. We believe that higher-dimensional data from in-vitro observations will lead to more sensitive calibrated parameters.

### v2

@@ Abstract
This study proposes a simple and practical approach based on in-vitro measurements and in-silico simulations using the likelihood-free Bayesian inference with the finite element method simultaneously for stochastic calibration and uncertainty quantification of the mechanical response of biological materials. We implement the approach for distal, middle, and proximal human Achilles tendon specimens obtained from diabetic patients post-amputation. A wide range of in-vitro loading conditions are considered, including one-step and two-step relaxation, as well as incremental cyclic loading tests. In-silico simulations are performed for the tendons assuming a fiber-reinforced viscoelastic response, which is modeled for the ground matrix and fiber components. Initially, the calibration of the specimen-specific parameters is predicted using Bayesian optimization and the sensitivity of each parameter is evaluated using the Sobol index and random forest. Then, these parameters are used as priors, and coupled with in-vitro data in simulation-based approximate Bayesian computation to calibrate and quantify the uncertainty parameters for three loading cases. The results demonstrate that in-silico simulations using the posterior parameters of approximate Bayesian computation, can capture the uncertainty bounds of in-vitro measurements. This approach provides a useful framework for stochastic calibration of constitutive material model parameters without the need to derive a likelihood function, regardless of the specimen’s geometry or loading conditions.
keywords:
In-vitro measurements, in-silico simulations, biological materials, uncertainty quantification, stochastic material calibration, approximate Bayesian computation, finite element method
journal: Methods
[…]
organization=Department of Mechanical Engineering, Ege University, addressline=35040, Bornova, city=Izmir, postcode=35040, country=Turkey
@@ 1 Introduction
In-silico mechanical modeling and simulation of biological materials have the potential to facilitate advancements in every aspect of the medical field. It offers numerous advantages, such as being a promising tool for the pre-clinical testing of medical devices, helping to diagnose and evaluate targeted treatments for patient-specific applications, and supplementing experimental investigations when in-vivo experiments are not possible [1, 2, 3, 4]. It can also model patient-specific tissue or organ-level behavior and, more generally, account for variations within a population. One such factor that significantly influences the in-silico simulation outcome is the constitutive material model, as well as the parameters used in it. The model and its calibrated biomechanical parameters should accurately reflect and mimic the tissue’s behavior. These parameters are typically calibrated using a non-linear least squares fit of the model to mechanical testing data from experiments such as tension [5, 6], compression [7], and biaxial testing [8]. For most calibration procedures, the choice of a constitutive model is assessed by the quality of fit (e.g., $R^{2}$ and the value of the root mean square error RMSE). This technique is efficient and powerful for patient or specimen-specific calibration. However, because most biological materials are nonhomogeneous and anisotropic, deploying experimental tests using standard mechanical tests may not be enough for mechanical characterization of the tissue. Therefore, calibration of the model parameters directly from the experimental data may not always be possible or relevant. In some studies, the inverse technique is performed to overcome these issues [9, 10]. In brief, the inverse approach is considered as the estimation of input parameters (material properties, geometric features, loading and boundary conditions, etc.) from output observations (displacement, strains, velocity). Once a forward model is created, the input parameters are predicted using experimental observations with an optimization technique. This method is powerful for patient/specimen-specific material calibration when inter-variability is not considered.
@@
In addition to inhomogeneous characteristics, biological materials often exhibit intrinsic variability in their biomechanical response when subjected to external loads. The inter-subject or inter-specimen variability and uncertainty in biomechanical properties of these materials are high, which may be influenced by many factors such as the donor’s age, genetics, the tissue’s pathological state, gender, and other specific factors related to the tissue characteristics. For in-vitro samples, embalming medium, temperature, humidity, and duration are the additional factors that can affect the biomechanical properties. Oftentimes, in-silico model is hindered not only by the incompleteness of the theoretical approach but also by uncertainties due to mentioned factors that arise in the experimental observations. Thus, there is an urgent need to ensure that it can capture the intrinsic population of variability and not only be representative of ‘mean’ individuals.
@@
Uncertainty can also arise due to lack of knowledge of the physical model of interest (insufficient initial and boundary conditions), as an addendum to the inherent variation in material properties [11, 12]. However, most of the studies implemented in the field of biomechanics have mainly investigated uncertainties due to variabilities in material properties using the Bayesian framework [13, 14]. In this framework, once the prior distribution of the parameters is constructed, in conjunction with a likelihood function, the posterior distribution is determined using random sampling methods. There has been much research in the field of brain mechanics [15, 16, 17, 18, 13]. Furthermore, several studies have established the utility of Bayesian inference in solid mechanics and soft tissues, addressing both parameter estimation and model selection. For instance, Madireddy et al. [15, 19] presented a Bayesian model selection and calibration framework for hyperelastic soft tissue models. Haughton et al. [20] used Bayesian techniques to quantify uncertainty in microstructural tendon models. Oden et al. [21] and Chiachío et al. [22] implemented Bayesian methods for selecting and validating phenomenological and damage progression models. Rappel et al. [23] provided detailed frameworks to identify the parameters of the elastoplastic material under uncertainty. In addition, James Beck and colleagues have applied Bayesian learning to structural health monitoring and inverse problems in mechanics  [24, 25, 26, 27]. Furthermore, Babuska et al. [28] presented a Bayesian model comparison approach for metallic fatigue data, while Xue et al. [29] proposed sparse Bayesian learning framework for robust damage localization using guided-wave testing. These studies primarily rely on standard Bayesian frameworks that require an explicit likelihood function. However, the likelihood function can be difficult or computationally expensive to derive for nonhomogeneous materials under complex biomechanical loading conditions. To overcome these such issues, an approximate Bayesian computation (ABC) approach that does not use the likelihood function has been proposed [30, 31] and implemented recently in some fields including structural health monitoring [32] medical decision making [33] and operation research management [34]. However, the coupling of in-silico simulations for stochastic calibration and uncertainty quantification using likelihood-free Bayesian computation in the field of mechanics of biological materials with high inter-specimen variability has been largely unaddressed to date, to the best of our knowledge. The most attractive aspect of this approach is that the likelihood function is not required to be explicitly represented in a closed analytical form. Once an in-silico model is built to mimic the in-vitro conditions, it directly bypasses the need for explicit likelihood evaluation and relies on comparing the simulated data from the in-silico model with experimental observations from the in-vitro test. This approach can be implemented in a wide range of in-vivo and in-vitro applications, ranging from predicting the mechanical characteristics change of tissue due to the risk of pathogenic disease to planning surgical procedures. We note that this study is not targeted to develop or present a constitutive material model for connective tissue, but rather to show how likelihood-free inference can be used directly with the numerical simulator and measurements. Any suggested constitutive model can be directly embedded in the simulator.
@@
In this work, we suggest a simple approach based on in-vitro measurements, in-silico simulations, and approximate Bayesian computation for stochastic calibration and uncertainty quantification of material parameters. The paper is structured as follows. First, a specimen-specific calibration procedure using Bayesian optimization is presented. Next, the sensitivity analysis and stochastic calibration steps implemented in ABC are detailed in the following section. Finally, the results obtained from the methodology are presented and summarized, along with some concluding remarks.
@@ 2.1 In-vitro measurements
The approach is performed on recently published data of Achilles tendons from diabetic patients who underwent below or above knee amputation [35]. In summary, the tendons were harvested from the proximal end to the insertion of the Achilles tendon at the calcaneus. Their length was approximately 120 to 130 cm and were partitioned into three equal sections (distal, middle, and proximal site). Then, they were resectioned to form a standard dog bone specimen. The dimensions of each specimen at the measurement site had a gauge length of approximately 10 mm, a width of $\sim$ 8 mm and a thickness of $\sim$ 4 mm. Uniaxial displacement-type loading was performed to collect force measurements. First, samples were preloaded at 0.1 MPa to achieve parallel alignment of the fibrils, and then preconditioned for 10 cycles with a strain amplitude of 2% strain at a strain rate of 0.5% $s^{-1}$. The test protocols were as follows:
  • 1.
@@
Further details of the tests and in-vitro measurements are reported in previous work [35]. A typical stress-strain observation of a specimen for three loading scenarios is presented in Figure 1. The force and displacement signals have a sampling rate of 100 Hz. In-vitro measurements include 81 datasets (9 specimens x 3 sites x 3 test protocols). Each of these data sets consists of three dimensions: the first is time, the second is displacement, and the third is force. The data sets are organized and saved as PYTHON pickle for further serialization processes between in-vitro measurements and in-silico simulations for sensitivity, optimization and Bayesian inference tasks.
@@ 2.2 Fiber reinforced visco elastic model
Soft biological tissues, such as the Achilles tendon, are collagenous tissues that connect the calf muscle to the calcaneus. Their primary function is to transmit tensile forces and act as mechanical dampers. These tissues are primarily composed of water, Type I collagen fibers, and proteoglycans. They can be viewed as fibrous composites, consisting of a ground matrix embedded with families of collagen fibers. The tendon is modeled using the fiber-reinforced viscoelastic (FRVE) approach, in which the solid matrix is divided into a nonfibrillar part representing the proteoglycan matrix and a fibrillar part representing the collagen fibers. The biomechanical response of these constituents, which drives the approximate Bayesian computation (ABC), is simulated using the FEBio framework [36]. In this FRVE model, two solid mixtures are considered: the ground matrix, representing the proteoglycan matrix, and the collagen fibrils, representing the fibers. Viscoelastic effects are incorporated for both constituents. Specifically, the ground matrix is modeled as a compressible visco linear elastic material, whereas the collagen fibrils are assumed to be compressible visco nonlinear elastic materials that respond only to tensile loads. The relaxation function for the fibril and matrix components is given by:
@@
where $t$ is time, $g_{i}$ and $\tau_{i}$ are $i$’th viscoelastic characteristic coefficients. The strain energy density function for the fiber component is defined as:
$\psi_{n}=\begin{cases}0,&I_{n}<1\quad\text{(under compression)}\\ \frac{\xi}{\alpha\beta}\left(\exp\left[\alpha(I_{n}-1)^{\beta}\right]-1\right),&1\leq I_{n}\leq I_{0}\quad\text{(nonlinear toe region)}\\ B(I_{n}-I_{0})-E_{f}\left(I_{n}^{1/2}-I_{0}^{1/2}\right)+\frac{\xi}{\alpha\beta}\left(\exp\left[\alpha(I_{0}-1)^{\beta}\right]-1\right),&I_{n}>I_{0}\quad\text{(linear region)}\end{cases}$ (2)
where $E_{f}$ is the fiber modulus in the linear region, $\alpha$ is the coefficient of the exponential term, and $\beta$ is the power-law exponent that controls the nonlinearity in the toe region. The scalar quantity $\lambda_{n}$ represents the fiber stretch ratio, defined as the ratio of the current fiber length to its reference length. The related invariant $I_{n}$ corresponds to the square of the fiber stretch ratio, $I_{n}=\lambda_{n}^{2}$. A threshold value $I_{0}=\lambda_{0}^{2}$ defines the transition from the nonlinear toe region to the linear region of the material response. The coefficients $\xi$ and $B$ are defined to ensure continuity of the strain energy function:
$\xi=\frac{E_{f}(I_{0}-1)^{2-\beta}\exp\left[-\alpha(I_{0}-1)^{\beta}\right]}{4I_{0}^{3/2}\left(\beta-1+\alpha\beta(I_{0}-1)^{\beta}\right)},$ (3)
$B=E_{f}\frac{2I_{0}\left(\beta-\frac{1}{2}+\alpha\beta(I_{0}-1)^{\beta}\right)-1}{4I_{0}^{3/2}\left(\beta-1+\alpha\beta(I_{0}-1)^{\beta}\right)}.$ (4)
In-vitro measurements showed that $\lambda_{0}$ is approximately 1.01 and assumed to be a fixed parameter. For this material type, the fiber modulus at the strain origin reduces to zero unless $\beta=2$. Thus, we set $\beta=2$ in the in-silico simulation. In summary, the matrix is assumed to exhibit viscoelastic behavior, while the fiber is modeled as a viscoelastic component that provides tensile reinforcement, using the fiber-exp-pow-linear material model available in FEBio. This approach captures both the time-dependent behavior and the load-bearing property of the fibers within the composite tissue. For additional theoretical background, see [36, 37]. The model parameters, their Bayesian optimization search ranges, and prior distributions are summarized in Table 1.
Table 1: Parameter space used in Bayesian optimization (BO) for the three sites.
@@ 2.3 In-silico simulations
Numerical methods such as the finite element method, meshless methods, or the finite volume method can be used as simulators to bridge the gap between in-vitro measurements and Bayesian computation for stochastic calibration and quantification of uncertainty in the biomechanical response of biological materials. Here, we used the finite element method as a simulator. The geometric dimensions of the in-vitro sample are approximately $10\times 8\times 4\,\text{mm}$ (gauge length $\times$ width $\times$ thickness). To reduce computational cost, the sample is modeled using a single 3D hexahedral element, with symmetry imposed in the y-direction by constraining the left face ($u_{y}(t)=0$), as shown in Figure 2. The bottom nodes are constrained vertically ($u_{z}(t)=0$), and the top nodes are displacement-controlled using measured displacements $u_{z}(t)$, consistent with the in-vitro test protocols. We note that additional in-silico outputs, such as the deformation gradient, the strain distribution, or the direct stress, can also be extracted and used with ABC if the corresponding measurements are available. Typically, full-field deformation measurements can be collected via image-based techniques, such as digital image correlation. We believe that higher-dimensional data from in-vitro observations will lead to more sensitive calibrated parameters.